OpenStudy (anonymous):

Can someone explain Induction to me? D:

OpenStudy (anonymous):

this is the way to show the subject of integers value always is right.

OpenStudy (anonymous):

Well, I know it's basically proofs.

OpenStudy (anonymous):

Suppose that P(n) is an issue related with integer number n. the problem is P(n) is right for all n>=m so you should show P(m) is corresct then assume that P(k) is correct then in conclusion show P(k+1) is correct.

OpenStudy (anonymous):

for example P(n)== 1+2+3....+n=n(n+1)/2 P(1) is 1=1*2/2=1 sup:1+2+3...+k=k(k+1)/2 so P(k+1) will be 1+2+3....+k+(k+1)=(k+1)(k+1+1)/2 it is enough that add k+1 in both sides of assume. and it'll show P(k+1) is true. that's it.